OpenStudy (anonymous):

You flip a coin three times. What is the probability that at least one of the outcomes will be tails?

OpenStudy (anonymous):

Would this be me multiplying 1/2 * 1/2 * 1/2 ? The thing that confuses me is when it says at least once.

OpenStudy (mathmale):

Have you considered drawing the set of all possible outcomes? Three independent tosses with 2 possible outcomes per toss ... ?

OpenStudy (anonymous):

I just picture flipping a coin threetimes. The probability of getting tails each time is 1/2

OpenStudy (faiqraees):

Make a probability tree which feature every outcome. Now find the probability of any outcome which have tails and then add every probability together.

OpenStudy (sparrow2):

in this problem there is no need of writing all: the opposite of at least one is 0 times,so first calculate the opposite this is 1/2*1/2*1/2=1/8.. now when you know the opposite ,for your statement 1-1/8=7/8

OpenStudy (sparrow2):

if you don't do this way,you should calculate tail 1 time,2 time,3 time and sum

OpenStudy (anonymous):

Okay, so the outcome would 7/8?

OpenStudy (sparrow2):

yeah

OpenStudy (anonymous):

Okay. Hmm. I understand that.

OpenStudy (sparrow2):

you can write every occurence but that takes long.this one is simpler and shorter

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Actually looking back at my notes, it does look like I have to subtract from one.

OpenStudy (mathstudent55):